OpenStudy (anonymous):

Solve: –4(2x – 1) = –6(x + 2) – 2

OpenStudy (anonymous):

-8x+4=-6x-12-2 (i distributed) -8x+4=-6x-14 (i simplified) -8x=-6x-18 (subtracted 4 from both sides to i can remove it from the left side) -2x=-16 (added 6x to both sides to comine the x's0) x=8 (divided both sides by -2)

OpenStudy (anonymous):

ok thanks